OpenStudy (anonymous):

A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a red marble, replacing it, and then picking a black marble?

OpenStudy (anonymous):

just to check my answer

OpenStudy (anonymous):

( 5/15)*(4/15)

OpenStudy (anonymous):

So this time, there are two difference. The marble colors are different for each pick and this problem has replacement, meaning you return the marble you picked out. In the beginning, the probability of a red marble is 5/15 After you've picked the first one and replace it, the probability of a black marble is 4/15 So your statement would be: Probability of red and black = Probability of 1st red * probability of 2nd black = (5/15) * (4/15) = 20/225 = 4/45

OpenStudy (anonymous):

thanks!!!!!!!!!! i got that answer